What is the center and the radius of the circle:
x^2+y^2-6y-16=0

Answer by lwsshak3(11628)

What is the center and the radius of the circle:
x^2+y^2-6y-16=0
***
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of the center, r=radius
x^2+y^2-6y-16=0
complete the square:
x^2+(y^2-6y+9)-9-16=0
x^2+(y-3)^2=25
center: (0,3)
radius=√25=5